OpenStudy (anonymous):

A system of equations is shown below: 6x-5y=5 3x+5y=4 The x-coordinate of the solution to this system of equations is:____ NUMERICAL ANSWERS EXPECTED!

OpenStudy (anonymous):

from first equations: x = (5 + 5y)/6 than you must substitute this expression into the second

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

3((5 + 5y)/6) + 5y = 4 you have to solve the equation

OpenStudy (anonymous):

y= 1/5...?

OpenStudy (anonymous):

i think 1/3

OpenStudy (anonymous):

but i'm not sure

OpenStudy (anonymous):

Okay, thanks so much! Could you help me with one more?

OpenStudy (anonymous):

give some time and i resolve it

OpenStudy (anonymous):

its not finish

OpenStudy (anonymous):

you must find x

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

How do I find x.?

OpenStudy (anonymous):

Do I just solve the equation orr...

OpenStudy (anonymous):

putting 1/5 in y's place?

OpenStudy (anonymous):

you must substitute y = 1/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay hold on!

OpenStudy (anonymous):

OKAY! X=38/15??